OpenStudy (anonymous):

4(1/2x+7)-3x=1/3(9-6x)+23

OpenStudy (anonymous):

thanks can you help me with-2n+9n-343=0

OpenStudy (anonymous):

sorry, noticed error in solution, retyping now

OpenStudy (anonymous):

(4/2x+28)-3x=(9-6x)/3+23

OpenStudy (anonymous):

2x+28-3x = 3-2x+23

OpenStudy (anonymous):

2x-3x+2x=26-28

OpenStudy (anonymous):

4x-3x=-2 x=-2

OpenStudy (anonymous):

I checked this solution also it is correct!

OpenStudy (anonymous):

are you sure the question isnt -2n^2 +9n-343=0?